Warnerwoods Company uses a periodic inventory system. It entered into the following purchases and sales transactions for March. Activities Units Sold at Retail Date Mar. 1 Beginning inventory Mar. 5 Purchase Mar. 9 Sales Mar. 18 Purchase Mar. 25 Purchase Mar. 29 Sales Units Acquired at Cost 145 units$80.00 per unit 445 units$85.00 per unit 465 units $115.00 per unit 210 units $90.00 per unit 290 units $92.00 per unit 250 units$125.00 per unit 715 units Totals 1,090 units For specific identification, the March 9 sale consisted of 90 units from beginning inventory and 375 units from the March 5 purchase; the March 29 sale consisted of 85 units from the March 18 purchase and 165 units from the March 25 purchase.
Answer #1

c) Average Cost method

Cost of goods available for sale : 145*80+445*85+210*90+290*92 = 95005

Total units : 1090 units

Average cost : 95005/1090 = 87.16

Cost of goods sold : 715*87.16 = 62319.79

Ending inventory : 375*87.16 = 32685.21

d) Specific identification

Cost of Googs sold : 90*80+375*85+85*90+165*92 = 61905

Ending inventroy = 55*80+70*85+125*90+125*92 = 33100